OpenStudy (anonymous):

y − 4 =3/4 (x + 8) , identify the equation of the same line in slope-intercept form

OpenStudy (cwrw238):

y − 4 =3/4 (x + 8) y = 3/4 (x + 8) + 4 y = 3/4 x + 6 + 4 y = 3/4 x + 10